Please help with:
Use substitution to determine whether -4, 2, and 1 are zeros of f(x)=x^3+3x^2-4x
Thanks.

Answer by themathprof(21)

f(-4)=(-4)^3+3(-4)^2-4(-4)
f(-4)=-64+3(16)+16
f(-4)=64+48+16
f(-4)=128 -4 is not a zero

f(2)=2^3+3(2)^2-4(2)
f(2)=8+3(4)-8
f(2)=8+12-8= 12 2 is not a zero

f(1)=1^3+3(1)^2-4(1)
f(1)=1+3(1)-4
f(1)=4-4=0 1 is a zero !!
visit www.math911.com.
download,install, activate math911 (all free!)
it has hundreds of problems like this worked
out step by step
tell your classmates and teacher